OpenStudy (anonymous):

The length of a rectangle is 2 more than 3 times the width. If the perimeter is 100 meters, what is the width of the rectangle? answer list 13 meters 12 meters 11 meters

terenzreignz (terenzreignz):

First step to an algebraic word problem? Represent the values :D If we let x = width of the rectangle What's the length?

OpenStudy (anonymous):

2+3x

terenzreignz (terenzreignz):

That's right :) The formula for the perimeter of a rectangle? P = ?

OpenStudy (anonymous):

s+s+s+s=p

terenzreignz (terenzreignz):

that's for a square :) The formula for the perimeter of a rectangle is 2L + 2W = P Now all you have to do is replace L, W, and P with the necessary values, and you're all set :)

OpenStudy (anonymous):

but how do i find the values

terenzreignz (terenzreignz):

I'll give you a hint: Your width (W) is x, right? so replace W with x: 2L + 2x = P Now you do L and P :)

OpenStudy (anonymous):

2(2+3x)+2x=100 4+6x+2x=100 8x+4=100 8x=96 x=12

OpenStudy (anonymous):

haha i had it right the first time

terenzreignz (terenzreignz):

You nailed it :D We were only looking for the width, right? so you got it:) What if we were looking for the length?

OpenStudy (anonymous):

i dont know how to that part

terenzreignz (terenzreignz):

You know your length is *two more than three times the width*, right? You already know your width, so...? :)

OpenStudy (anonymous):

u get rid of of width

terenzreignz (terenzreignz):

What do you mean by that?

OpenStudy (anonymous):

no divid width by 3 then times by 5 wat was i thinking

terenzreignz (terenzreignz):

Ok, I'll be blunt this time around ;) x is our width our length is 2 + 3x and recently, we found x = 12 ..., so our length, which is (2 + 3x) must be...? (Hint: Substitute 12 for x) :D

OpenStudy (anonymous):

(2+3x)=38

terenzreignz (terenzreignz):

That's right :) Nicely done. A bit more practice and you should be breezing through these algebraic word problems :)

OpenStudy (anonymous):

i hope an thx alot again

terenzreignz (terenzreignz):

No problem, and keep it up :)

OpenStudy (anonymous):

ok i will